OpenStudy (anonymous):

Can anyone help me set this word problem up please? A minor league baseball team plays 97 games in a season. if the team won 13 more than three times as many games as they lost, how many wins and loses did the team have?

OpenStudy (auctoratrox):

let x represent the number of losses. so now the # of games won is represented by 3x+13, now add them both and set equal to 97. so we get 3x+13+x=97. do you get how I did that?

OpenStudy (lgbasallote):

let w = number of winds l = number of lose it says they played 97 games so w + l = 97 it also says that the team won 13 more than 3 times the number of games they lost so win = 13 + 3 x lose w = 13 + 3x

OpenStudy (anonymous):

yes, that makes sense. thank you